A model rocket is launched from a platform and its height h metres is given as a function of the elapsed time t seconds by h=-4.9t*2+180t+2. Determine the following:
a)The maximum height of the rocket
b)The time required for the rocket to reach its maximum height
c)The time required for the rocket to reach the ground
Thanks!

Answer by Alan3354(69443)

h=-4.9t^2 + 180t + 2.
Determine the following:
b)The time required for the rocket to reach its maximum height
max ht is the vertex, at t = -b/2a
t = -180/9.8 =~ 18.367 seconds
-------------
a)The maximum height of the rocket
Sub 18.367 for t
Max ht = -4.9*18.367^2 + 180*18.367 + 2
==================
c)The time required for the rocket to reach the ground
-4.9t^2 + 180t + 2 = 0
Solve for t.